# Generalizing Verifiable Instruction Following

### v2

@@ 6 Related Work
Following instructions precisely and adhering to specific output or formatting constraints is often still a challenge for language models. LLMs’ reasoning abilities, for example, decline when they also have to adhere to formatting constraints tam2024let . And models can struggle at doing constrained generation with fine-grained constraints sun2023evaluating . Previous works improved models’ instruction following abilities by scaling up the instruction fine-tuning stage chung2024scaling , through activation steering stolfo2024improving , DPO training kim-etal-2025-systematic , and using RL with verifiable rewards lambert2024t .
@@ Appendix A Out-of-Distribution Test Constraints
count person_names Mention at least {N} different person names in the response.
@@
count words_french Every {N}th word of your response must be in french.
@@
[…]
@@
ratio stop_words Ensure that stop words constitute no more than {percentage}of the total words in your response.
@@
copy repeat_span Copy the span of words that lies between (and including) index n_start and n_end, the indices are word indices, split by whitespace!
@@ Appendix C LLM-as-judge Prompt
Appendix D Chat Template

### v3

@@ 6 Related Work
Following instructions precisely and adhering to specific output or formatting constraints is often still a challenge for language models. LLMs’ reasoning abilities, for example, decline when they also have to adhere to formatting constraints tam2024let. And models can struggle at doing constrained generation with fine-grained constraints sun2023evaluating. Previous works improved models’ instruction following abilities by scaling up the instruction fine-tuning stage chung2024scaling, through activation steering stolfo2024improving, DPO training kim-etal-2025-systematic, and using RL with verifiable rewards lambert2024t. Reinforcement learning with verifiable rewards has been shown to be an effective approach for improving mathematical reasoning and coding abilities in language models lambert2024t; shao2024deepseekmath; guo2025deepseek; zelikman2022star; hoffman2023training; kazemnejad2024vineppo, with theoretical support for its viability amit2024models.
@@ Appendix A Out-of-Distribution Test Constraints
count person_names Mention at least {N} different person names in the response, from this list of person names: Emma, Liam, Sophia….
@@
count words_japanese Every {N}th word of your response must be in Japanese.
@@
ratio overlap Maintain a trigram overlap of {percentage}% ($\pm$2%) with the provided reference text.
@@
ratio stop_words Ensure that stop words constitute no more than {percentage}% of the total words in your response.
@@
copy repeat_span Copy the span of words that lies between (and including) index n_start and n_end, the indices are character indices!
@@ Appendix C LLM-as-judge Prompt
Appendix D Analysis
D.1 Length Analysis
The single-turn IFBench prompts have on average a length of 76 tokens and the multi-turn conversations provide an average input length of 408 tokens (consisting of prompt, response, prompt). The frontier models on the left side of Figure 1 generate IFBench responses that are on average 2214 token long, and the IF-RLVR trained models tend to generate shorter reponses, with an average token length of 210.
D.2 Constraint Diversity
Compared to IFEval [zhou2023instruction], which covers atomic constraints focusing on keyword frequency, basic formatting, template-based responses and simple transformations, IFBench’s constraints are a bit more challenging. The more complex constraints include mathematical or algorithmic requirements, such as trigram overlap percentages and alphabet cycling, and nested linguistic patterns, such as palindromes and alliteration progression. Besides these more adversarial constraints, we also included atomic constraints similar but different to IFEval: i.e. “Include exactly {N} numbers in the response.", where IFEval asks to include certain keywords, but does not specify how many times and that they should be numbers. As all constraints were human-written, we could ensure that there is no overlap between the two evaluation sets.
Appendix E Mitigating Reward Hacking
While GRPO training with verifiable rewards for precise IF is great at teaching LLMs to follow output constraints, it can sometimes result in models that over-prioritize the constraint over the full instruction. An example of such an output is given in Figure 7. This could also be called over-optimization. We propose adding a general reward model (RM) signal to the verifiable reward. The intuition is that while the verifiable reward checks for the adherence to the output constraint, the general reward model provides signal for whether the response answers the prompt. We apply the reward from the reward model to all generations that received a verifiable reward >0>0, as follows: For a batch of data, we assign the final reward, $F_{i}$ to that instance corresponding to conditions of the verifiable reward, $V_{i}$, and the reward model score, $S_{i}$:
We use the openly available Llama-3.1-Tulu-3-8B-RM as our general preference reward model, set $\alpha=7$ [^4] We chose 7 as this is around the mean reward score given by the model over a large set of instances., and use an effective batch size of 512, with 8 samples/prompt.
Appendix F Chat Template
@@
Appendix G Evaluation Details
When we generate for evaluation, we generally use a temperature of 0 and adjust the maximum generated tokens depending on the model type, i.e. for thinking models we allow to generate more tokens and we then process the output to extract the answer without the reasoning chains. When evaluating Deepseek R1, we used their recommended generation settings of temperature = 0.6 and top p = 0.95, and for o3 we use their required temperature of 1.